OpenStudy (odys):

(-1/2)(4*)(-2)(7y)(-1) please help will fan and medal!

OpenStudy (odys):

A -28xy B -28 C 28xy D 27

OpenStudy (odys):

(-1/2)(4*)(-2)(7y)(-1 simplify A -28xy B -28 C 28xy D 27

OpenStudy (anonymous):

Can you write your question more clearly?

OpenStudy (anonymous):

(4*) = ??

OpenStudy (odys):

yeah hold on first time asking question

OpenStudy (anonymous):

First time? Really? I don't think so..

OpenStudy (odys):

i answer questions not ask most the time

OpenStudy (odys):

and i answer science

OpenStudy (anonymous):

That is great..

OpenStudy (anonymous):

you can use Draw Tool.. Can't you?

OpenStudy (odys):

yeah but like to complicated right now i'll try to explain better

OpenStudy (odys):

okay simplify (-1/2)(4*)(-2)(7y)(-1)

OpenStudy (odys):

answers are a -28xy B -28 C 28xy D 27

OpenStudy (anonymous):

I am writing nicely for you, check I am right or not..

OpenStudy (odys):

k thnx

OpenStudy (anonymous):

\[\frac{-1}{2} \times 4 \times (-2) \times 7y \times (-1)\]

OpenStudy (odys):

and then solve this then or?

OpenStudy (odys):

do i solve this

OpenStudy (anonymous):

yes..